An angle is 14 degrees more than 4 times its' supplement. Find the angle and its' supplement.
Answer by checkley77(12569)

x+y=180
x=4y+14
4y+14+y=180
5y=180-14
5y=166
y=166/5
y=33.2 degrees.
x+33.2=180
x=180-33.2
x=146.8 degrees.
Proof:
146.8=4*33.2+14
146.8=132.8+14
146.8=146.8